# TransDiffuser: End-to-end Trajectory Generation with Decorrelated Multi-modal Representation for Autonomous Driving

### v1

@@ Abstract
In recent years, diffusion model has shown its potential across diverse domains from vision generation to language modeling. Transferring its capabilities to modern autonomous driving systems has also emerged as a promising direction. In this work, we propose TransDiffuser, an encoder-decoder based generative trajectory planning model for end-to-end autonomous driving. The encoded scene information serves as the multi-modal conditional input of the denoising decoder. To tackle the mode collapse dilemma in generating high-quality diverse trajectories, we introduce a simple yet effective multi-modal representation decorrelation optimization mechanism during the training process. TransDiffuser achieves PDMS of 94.85 on the NAVSIM benchmark, surpassing previous state-of-the-art methods without any anchor-based prior trajectories.
Keywords: End-to-end Autonomous Driving, Representation Learning
1 Introduction
@@
Autonomous driving [1, 2] has gained widespread attention from both academia and industry for its potential in improving traffic safety and efficiency. Substantial advancements have been attained across a diverse range of tasks, including real-time localization [3, 4], scene perception [5, 2] and motion planning [6]. Early motion planning methods typically involve a sequential process of perception, prediction, and subsequent planning, which leads information loss and cascading latency [7]. In recent years, an important research direction is to develop planning-oriented autonomous driving systems [1] in a fully end-to-end manner. With raw sensor data as input, an end-to-end driving model is expected to directly output a desirable trajectory to guide future motion planning. Early works [1, 8] aim to produce a single plausible trajectory by merely imitating the annotated human expert driving behaviors in the training dataset.
@@
However, this data-driven paradigm may not be fully capable to generalize to the real-world scenarios. Considering the existence of diverse yet complex driving scenarios and different feasible driving styles [9], there is rarely only one feasible trajectory [10]. Recent attempts [11, 12, 13, 10] have increasingly focused on generating multi-mode future trajectories as feasible candidates [^1] In this paper, multi-mode trajectories denote generating multiple possibly feasible candidate trajectories.. To generate multi-mode high-quality trajectories from the continuous action space, one research line, represented by Hydra-MDP series [13, 14, 15], simplifies this challenge into selecting most feasible candidates from a fixed planning vocabulary to discretize the action space. Another notable research trend [10, 16, 9] aims to transfer the success of diffusion models [17] in robotics domain [18] to generate multi-mode trajectories in end-to-end autonomous driving domain, using scene and motion information as conditional input to produce multi-mode trajectories as candidates. Similar to Hydra-MDP series, GoalFlow [10] imposes a strong constraint on the trajectory generation process by establishing a dense vocabulary of anchor trajectory endpoints. DiffusionDrive [16] further highlights the issue of mode collapse, wherein the generated trajectories lack diversity, as different random noise inputs tend to converge to similar trajectories during the denoising process. Therefore, it implements a truncating diffusion policy by partitioning the Gaussian distribution into multiple sub-Gaussian distributions centered around prior anchor trajectories.
@@
Following the intuition of the above recent Diffusion based approach, we propose TransDiffuser, an encoder-decoder based multi-mode trajectory generation model. We utilize the frozen Transfuser backbone [8] to encode the scene perception information. Then, the scene information from front-viewed cameras and LiDAR, and motion information of the current ego vehicle is encoded as the conditional input of the Diffusion based denoising decoder. Unlike previous works, we identify another underlying bottleneck that leads to mode collapse in generated trajectories: The under-utilization of the encoded multi-modal conditional information. To generate more diverse feasible trajectories, we introduce a computation-efficient yet effective plug-and-play multi-modal representation decorrelation optimization mechanism during the denoising process, which aims to better exploit the multi-modal representation space to guide more diverse feasible planning trajectories from the continuous action space. As a result, the proposed method achieves the new state-of-the-art performance on the planning-oriented NAVSIM benchmark without any explicit guidance like anchor-based trajectories. To sum up, our contributions can be summarized as follows:
@@
  • We propose an encoder-decoder generative trajectory model TransDiffuser. It firstly encode the scene perception and motion of the ego vehicle, and then utilize the encoded information as conditional input of Denoising Decoder to decode multi-mode diverse feasible trajectories.
  • To further increase the diversity of generated trajectories, we introduce a computation-efficient multi-modal representation decorrelation mechanism during the training process.
  • We achieve the new state-of-the-art PDM score 94.85 on the NAVSIM benchmark without any explicit guidance like anchor trajectories or predefined vocabulary.
2 Related Works
2.1 End-to-end Autonomous Driving
@@
Modern autonomous driving systems [1, 16, 12, 13, 9] increasingly adopt end-to-end learning paradigms to directly map raw sensor data to driving decisions, streamlining the process from perception to action. Herein we review most recent advances in end-to-end autonomous driving for motion planning task. We categorize current methods into three groups as shown in Table 1.
@@
Auto-regressive Based Models. This approach is relatively traditional, which typically predict a trajectory via sequential Auto Regressive (AR) generating. To our best knowledge, UniAD [1] is the pioneering work showcasing the potential of end-to-end autonomous driving by integrating multiple perception tasks to facilitate planning. Transfuser [8] integrates image and LiDAR representations using self-attention mechanism and uses GRU to yield the planning trajectory in an auto-regressive paradigm. Its variant, LTF, is a light-weight image-only version where the LiDAR backbone is replaced by a learnable embedding. PARA-Drive [19] proposes a computationally efficient system which performs mapping, planning, motion prediction, and occupancy prediction tasks in parallel.
@@
Scoring Based Models. This approach often generates multi-mode trajectories and selects optimal trajectory with the designed scoring functions. VADv2 [12] is an early attempt to perform multi-mode planning by scoring and sampling from a large fixed vocabulary of anchor trajectories. Centaur [20] employs cluster entropy to measure uncertainty by analyzing the distribution of multiple trajectory candidates generated by the planner. WoTE [21] leverages a latent BEV world model to forecast future BEV states for trajectory evaluation. Hydra-MDP series [13, 14, 15] convert the trajectory generation task into selecting optimal trajectory from predefined trajectory vocabulary. and propose an expert-guided hydra distillation strategy to align the planner with simulation-based metrics.
Diffusion Based Models. GoalFlow [10] employs flow matching to generate goal-conditioned, multi-mode trajectories and introduces a trajectory scorer to efficiently select optimal trajectory using the goal point as a reference. DiffusionDrive [16] uses truncated diffusion policy to address the issues of mode collapse and reduce heavy computational overhead by denosing from anchored Gaussian distributions ($\dagger$ in Table 1) instead of anchor trajectories with fewer steps while preserving trajectory diversity. TrajHF [9] pioneers to introduce additional reinforcement learning on their private dataset to fine-tune the generative trajectory model.
@@
Table 1: Comparison of existing models. * denotes generating multi-mode candidate trajectories.
Model Venue Year Image Encoder LiDAR Anchor Paradigm
Transfuser [8] ICRA & TPAMI 2023 ResNet-34 $\checkmark$ AR
UniAD [1] CVPR 2023 ResNet-34 AR
PARA-Drive [19] CVPR 2024 ResNet-34 AR
VADv2 [12] 2024 ResNet-34 $\checkmark$ Scoring*
Hydra-MDP [13] CVPRW 2024 ResNet-34 $\checkmark$ $\checkmark$ Scoring*
[…]
GoalFlow [10] CVPR 2025 ResNet-34 $\checkmark$ $\checkmark$ Diffusion*
Hydra-MDP++ [14] 2025 ResNet-34 $\checkmark$ Scoring*
Hydra-NeXt [15] 2025 ResNet-34&V2-99 $\checkmark$ Scoring*
TrajHF [9] 2025 ViT $\checkmark$ Diffusion*
Centaur [20] 2025 ResNet-34 $\checkmark$ $\checkmark$ Scoring*
WoTE [21] 2025 ResNet-34 $\checkmark$ $\checkmark$ Scoring*
TransDiffuser (Ours) 2025 ResNet-34 $\checkmark$ Diffusion*
2.2 Generative Trajectory Model
@@
Generative models have garnered significant attention in both academic and industry due to their remarkable capabilities in producing realistic data across various domains. These models have been successfully applied in numerous tasks, such as image synthesis using Generative Adversarial Network [22] and Variational Autoencoder [23], text generation with large language models [24], and audio synthesis with WaveNet [25]. Diffusion models have been proposed and widely studied in recent years, and demonstrated their powerful generative abilities across vision generation [17, 26, 27, 28] and language generation [29]. A recent shift has emerged in the field of end-to-end autonomous driving, where researchers [16, 10, 9] embark on exploring the use of Diffusion models for planning tasks. Diffusion models, known for their ability to generate high-quality data by iteratively denoising random noise, have shown promise in generating smooth and realistic driving trajectories. However, pointed by DiffusionDrive [16], one bottleneck for diffusion models is that diffusion based trajectory generative models showcases the less diversity of the decoded denoised trajectories. This issue is often called by mode collapse. It proposes truncated diffusion policy that begins the denoising process from an anchored gaussian distribution instead of a standard Gaussian distribution. GoalFlow [10] used a goal point vocabulary to mitigate the mode collapse issue to assist the efficient trajectory genration. Different from above works, we aim to mitigate the mode collapse issue via improving the fused intermediate multi-modal representations, which assists to increase the diversity of final decoded trajectories. Notably, we do not utilize any form of guide information such as anchors or vocabulary, allowing for more flexible and diverse planning.
2.3 Representation Learning
@@
Representation learning has shown its potential in many deep learning applications like self-supervised learning [30], supervised classification [31, 32, 33], noisy label learning [34, 35, 36] and reinforcement learning [37]. In general, these attempts aim to fully exploit the representation space to learn better intermediate representations for downstream applications. One popular approach is contrastive learning [38, 39] by mining the representation similarities among positive and negative examples. However, this approach generally requires very large training batch (e.g. 4096 or 8192) to fully extract reliable supervision from batch examples, which may be not suitable and computation-efficient for training autonomous driving models. Inspired by recent representation learning works [37, 34, 31, 40], we optimize the intermediate representations in our task by decoupling the interactions of different representation dimensions. This optimization approach does not rely on the large training batchsize and bring little computation overhead as analyzed in [37, 31], which we will introduce in Section 3.4.
3 Method
3.1 Preliminary
@@
Figure 1: Overview of the proposed TransDiffuser architecture. Note that the parameters of the scene perception (images and LiDAR) encoder are frozen.
3.2 Scene & Motion Encoder
@@
3.3 Denoising Decoder
The Denoising Decoder (Figure 1(b)) is responsible for generating the planned trajectory by iteratively refining an initial noise estimate, conditioned on the encoded scene and motion features derived from the Scene Encoder. It takes the encoded feature group $feat=\{F_{bev},F_{img},F_{LiDAR},Emb_{action},Emb_{ego}\}$ as conditional input. Different features within this group are sequentially fused via multi-head cross-attention [41]. The decoder then outputs a feasible action from the continuous action space. By periodically accumulating these decoded actions, the final trajectory is formed.
To achieve this generation and handle the complexity of driving scenarios, we employ Denoising Diffusion Probabilistic Models (DDPM) [17] as the optimization framework. DDPM involves forward process where noise is gradually added to the ground truth data, and reverse process where the model learns to denoise. Following [9, 16], we focus on reverse denoising process, which transitions from Gaussian noise towards noise-free state $x_{0}$. This transition is governed via this equation:
@@ 3.3 Denoising Decoder
where $z\sim\mathcal{N}(0,I)$, $t\in\{1,...,\mathbf{T}\}$ denotes the noise level timestep, $\mathbf{T}$ is the total number of denoising steps, and $\epsilon_{\theta}$ represents our denoising decoder parameterized by $\theta$. The noise prediction $\epsilon_{\theta}$ is conditioned on the noisy state $x_{t}$, the timestep $t$, and the feature group $feat$. The parameters $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\sigma_{t}^{2}=\beta_{t}$ are derived from a predefined noise schedule $\beta_{t}$. Essentially, Equation 2 describes how to estimate a slightly less noisy state ($x_{t-1}$) from the current noisy state ($x_{t}$) using the model’s noise prediction ($\epsilon_{\theta}$).
The decoder is trained to accurately predict the noise added during the forward process. The optimization objective involves minimizing the difference between the actual sampled noise $\epsilon$ and the predicted noise $\epsilon_{\theta}$, formulated as a gradient descent step on:
$\nabla_{\theta}||\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t,feat)||^{2}$ (3)
where $\epsilon$ is randomly sampled from $\mathcal{N}(0,I)$. The detailed computation of this trajectory denoising loss, $\mathcal{L}_{diff}$, using Mean Squared Error (MSE) is outlined in Algorithm 1.
Input: Trajectory GT $label$, Scene feature $feat$, Number of denoising timesteps $\mathbf{T}$
@@
1 $noise\leftarrow\text{Gaussian}(label.shape)$ ; // Generate random Gaussian noise
2 $step\leftarrow\text{Rand}(0,\mathbf{T})$ ; // Pick a random time step
3 $noisy\_target\leftarrow\text{AddNoise}(label,noise,step)$ ; // Apply noise in the forward process
4 $pred\leftarrow\text{Denoising\_Decoder}(noisy\_target,step,feat)$ ; // Denoising process

### v2

# TransDiffuser: Diverse Trajectory Generation with Decorrelated Multi-modal Representation for End-to-end Autonomous Driving

@@ Abstract
In recent years, diffusion models have demonstrated remarkable potential across diverse domains, from vision generation to language modeling. Transferring its generative capabilities to modern end-to-end autonomous driving systems has also emerged as a promising direction. However, existing diffusion-based trajectory generative models often exhibit mode collapse where different random noises converge to similar trajectories after the denoising process. Therefore, state-of-the-art models often rely on anchored trajectories from pre-defined trajectory vocabulary or scene priors in the training set to mitigate collapse and enrich the diversity of generated trajectories, but such inductive bias are not available in real-world deployment, which can be challenged when generalizing to unseen scenarios. In this work, we investigate the possibility of effectively tackling the mode collapse challenge without the assumption of predefined trajectory vocabulary or pre-computed scene priors. Specifically, we propose TransDiffuser, an encoder-decoder based generative trajectory planning model, where the encoded scene information and motion states serve as the multi-modal conditional input of the denoising decoder. Different from existing approaches, we exploit a simple yet effective multi-modal representation decorrelation optimization mechanism during the denoising process to enrich the latent representation space which better guides the downstream generation. Without any pre-defined trajectory anchors or pre-computed scene priors, TransDiffuser achieves the PDMS of 94.85 on the closed-loop planning-oriented benchmark NAVSIM, surpassing previous state-of-the-art methods. Qualitative evaluation further showcases TransDiffuser generates more diverse and plausible trajectories which explore more drivable area.
I Introduction
@@
Recently, substantial advancements have been attained across a diverse range of autonomous driving tasks including real-time localization [1, 2] and scene perception [3, 4]. Among them, planning-oriented autonomous driving [4, 5] has gained widespread attention from both academia and industry for its potential in improving traffic safety and efficiency. Early planning approach usually adopts a sequential process of perception, prediction, and subsequent planning, which causes information loss and cascading latency [6]. Over the past few years, developing fully end-to-end planning-oriented autonomous driving systems [5] has emerged as a key research direction. Taking raw sensor data as input, an end-to-end driving model is expected to directly output an optimal trajectory for guiding future motion planning. Early works [5, 7] aim to generate a single plausible trajectory in an auto-regressive manner by simply imitating annotated human expert driving behaviors in the training dataset.
@@
However, this data-driven paradigm can be not fully capable to generalize to unseen and real-world scenarios. Considering the existence of diverse yet complex driving scenarios and different feasible driving styles [8], there is rarely single feasible trajectory [9]. Recent attempts [10, 11, 9] have increasingly focused on generating multi-mode trajectories as feasible candidates [^1] In this paper, multi-mode trajectories denote generating multiple possibly feasible candidate trajectories.. To generate multi-mode trajectories from the continuous action space, one research line, represented by Hydra-MDP series [11, 12, 13], simplifies this challenge into selecting feasible candidates from a fixed planning vocabulary to discretize the action space. Another emerging research line [9, 14, 8] aims to transfer the success of diffusion models [15, 16] to generate multi-mode trajectories, using scene and motion information as conditional input to produce multi-mode trajectories as candidates. GoalFlow [9] imposes a constraint on the trajectory generation process with scene priors, which is achieved by establishing a dense vocabulary of goal points. DiffusionDrive [14] further highlights the challenge of mode collapse, wherein generated trajectories lack diversity as different noise inputs tend to converge to similar trajectory distribution after denoising. It partitions the Gaussian distribution into multiple sub-Gaussian distributions centered around prior anchor trajectories for initialization. Notably, both previous research lines often require the pre-definition of trajectory vocabulary for anchor trajectories or pre-computation of scene priors. This introduces inductive bias which can face challenges for unseen scenarios.
@@
Following the intuition of the above Diffusion based approach, we propose TransDiffuser, an encoder-decoder based multi-mode trajectory generation model. We utilize the frozen Transfuser backbone [7] to encode the scene perception information from front-viewed cameras and LiDAR. The scene information and motion information of the current ego vehicle is then encoded as the conditional input of the Diffusion based denoising decoder. Unlike previous works, we identify another underlying bottleneck that leads to mode collapse in generated trajectories: The under-utilization of the encoded multi-modal representation from the conditional input of different modalities. To explore the possibility of generating diverse and feasible trajectories without any anchors or scene priors, inspired by recent advance in self-supervised representation learning, we exploit a computation-efficient yet effective plug-and-play multi-modal representation decorrelation optimization mechanism during the denoising process, which aims to better exploit the multi-modal representation space to guide more diverse feasible planning trajectories from the continuous action space. TransDiffuser achieves the new state-of-the-art performance on the planning-oriented NAVSIM benchmark without any explicit guidance like anchor-based trajectories or scene priors, and qualitative analysis further confirms it can generate diverse trajectories which explore more drivable space. To sum up, our contributions are outlined as follows:
@@
  • We propose an encoder-decoder based generative trajectory model TransDiffuser. It firstly encodes the scene perception and motion state of the ego vehicle, and then utilizes the encoded information as conditional input of denoising decoder to decode multi-mode diverse yet feasible trajectories.
  • Different from existing works that rely on predefined trajectories or pre-computed scene priors, we exploit a computation-efficient multi-modal representation decorrelation mechanism during the denoising process to enhance the diversity of generated trajectories to address the model collapse dilemma.
  • We achieve the new state-of-the-art PDMS 94.85 on NAVSIM benchmark without any explicit guidance like predefined anchor trajectories or scene priors. Qualitative analysis showcases TransDiffuser generates diverse yet feasible trajectories which better explore the drivable space.
II Related Works
II-A End-to-end Autonomous Driving
@@
Modern autonomous driving systems [5, 14, 11, 8] increasingly adopt end-to-end learning paradigms to directly map raw sensor data to driving decisions, streamlining the process from perception to action. Herein we review most recent related advances for planning and categorize current methods into three groups as shown in Table I.
Figure 1: Main approaches for end-to-end autonomous driving.
@@
Auto-regressive Based Models. This approach is relatively traditional, which typically predict a trajectory via sequential auto-regressive (AR) generating. To our best knowledge, UniAD [5] is the pioneering work showcasing the potential of end-to-end autonomous driving by integrating multiple perception tasks to facilitate planning. Transfuser [7] integrates image and LiDAR representations using self-attention and uses GRU to yield the planning trajectory in an auto-regressive paradigm. Its variant LTF is a light-weight image-only version where the LiDAR backbone is replaced by a learnable embedding. PARA-Drive [17] proposes a computation-efficient system which performs mapping, planning, motion prediction and occupancy prediction tasks in parallel.
@@
Scoring Based Models. This approach often generates multi-mode trajectories and selects optimal trajectory with the designed scoring functions. VADv2 [18] is an early attempt to perform multi-mode planning by scoring and sampling from a large fixed vocabulary of anchor trajectories. Centaur [19] employs cluster entropy to measure uncertainty by analyzing the distribution of multiple trajectory candidates generated by the planner. WoTE [20] leverages a latent BEV world model to forecast future BEV states for trajectory evaluation. Hydra-MDP series [11, 12, 13] convert the trajectory generation task into selecting optimal trajectory from predefined trajectory vocabulary. and propose an expert-guided hydra distillation strategy to align the planner with simulation-based metrics. R2SE [21] introduces a reinforced refinement framework with 3D backbone BEVFormer [22] that improves scaled hard case performance.
Diffusion Based Models. One emerging trend aims to transfer the generative capabilities of difffusion models into the end-to-end planning task. Early attempts include GoalFlow [9], DiffusionDrive [14] and TrajHF [8], which will be detailedly discussed in the later Section II-B.
@@
TABLE I: Comparison of existing models. * denotes generating multi-mode candidate trajectories.
Model Venue Year LiDAR Anchor Paradigm
Transfuser [7] ICRA & TPAMI 2023 $\checkmark$ AR
UniAD [5] CVPR 2023 AR
PARA-Drive [17] CVPR 2024 AR
VADv2 [18] 2024 $\checkmark$ Scoring*
Hydra-MDP [11] CVPRW 2024 $\checkmark$ $\checkmark$ Scoring*
DiffusionDrive [14] CVPR 2025 $\checkmark$ $\checkmark$ Diffusion*
GoalFlow [9] CVPR 2025 $\checkmark$ $\checkmark$ Diffusion*
Hydra-MDP++ [12] 2025 $\checkmark$ Scoring*
Hydra-NeXt [13] 2025 $\checkmark$ Scoring*
R2SE [21] 2025 Scoring*
TrajHF [8] 2025 $\checkmark$ Diffusion*
DIVER [23] 2025 $\checkmark$ Diffusion*
Centaur [19] 2025 $\checkmark$ $\checkmark$ Scoring*
WoTE [20] ICLR 2025 $\checkmark$ $\checkmark$ Scoring*
TransDiffuser 2025 $\checkmark$ Diffusion*
II-B Generative Trajectory Model
@@
Generative models have garnered significant attention due to their remarkable capabilities in producing realistic data across various domains. These models have been successfully applied in numerous tasks, such as image synthesis using Generative Adversarial Network [24] and Variational Autoencoder [25] and text generation with large language models [26]. Diffusion models have been proposed and widely studied in recent years, and demonstrated their powerful generative abilities across vision generation [15, 28, 29, 30] and language generation [31]. A recent shift has emerged in the field of end-to-end autonomous driving, where researchers [14, 9, 8] embark on exploring Diffusion models for planning. Diffusion models, known for their ability to generate high-quality data by iteratively denoising random noise, have shown promise in generating smooth and realistic driving trajectories. However, pointed by DiffusionDrive [14], one bottleneck lies in that diffusion based trajectory generative models showcases the less diversity of the decoded denoised trajectories. This issue is often called by mode collapse. It proposes truncated diffusion policy that begins the denoising process from an anchored gaussian distribution instead of a standard Gaussian distribution. GoalFlow [9] used a goal point vocabulary to mitigate the mode collapse issue to assist the efficient trajectory generation. TrajHF [8] pioneers to introduce preference optimization on their private dataset to align the generative trajectory model. Different from above works, we aim to mitigate the mode collapse issue via improving the fused intermediate multi-modal representations, which assists to increase the diversity of final decoded trajectories. Notably, we do not utilize any form of guidance like anchors or scene priors, allowing for more flexible and diverse planning.
II-C Representation Learning
@@
Representation learning has shown its potential in many deep learning applications like self-supervised learning [32], supervised classification [33, 34, 35], noisy label learning [36, 37, 38] and reinforcement learning [39]. In general, these attempts aim to fully exploit the representation space to learn better intermediate representations for downstream applications. One popular approach is contrastive learning [40] by mining the representation similarities among positive and negative examples. However, this approach generally requires very large training batch (e.g. 4096 or 8192) to fully extract reliable supervision from batch examples, which may be not suitable and computation-efficient for training autonomous driving models. Inspired by the recent advance [39, 36, 33], we optimize the intermediate representations in our task by decoupling the interactions of different representation dimensions. This optimization approach does not rely on the large training batchsize and bring little computation overhead as analyzed in [39, 33], which we will introduce in Section III-D and IV-C.
III Method
III-A Preliminary
@@
Figure 2: Overview of the proposed TransDiffuser architecture. We freeze the parameters of the scene encoder.
III-B Scene & Motion Encoder
@@
III-C Denoising Decoder
The denoising decoder (Figure 2(b)) is responsible for generating the planned trajectory by iteratively refining an initial noise estimate, conditioned on the encoded scene and motion features derived from the Scene Encoder. It takes the encoded feature group $feat=\{F_{bev},F_{img},F_{LiDAR},Emb_{action},Emb_{ego}\}$ as conditional input. Different features within this group are sequentially fused via multi-head cross-attention [42]. The decoder then outputs a feasible action from the continuous action space. By periodically accumulating these decoded actions, the final trajectory is formed. To achieve this generation and handle the complexity of driving scenarios, we employ Denoising Diffusion Probabilistic Models (DDPM) [15] as the optimization framework. DDPM involves forward process where noise is gradually added to the ground truth data, and reverse process where the model learns to denoise. Following [8, 14], we focus on reverse denoising process, which transitions from Gaussian noise towards noise-free state $x_{0}$. This transition is governed via this equation:
@@ III-C Denoising Decoder
where $z\sim\mathcal{N}(0,I)$, $t\in\{1,...,\mathbf{T}\}$ denotes the noise level timestep, $\mathbf{T}$ is the total number of denoising steps, and $\epsilon_{\theta}$ represents our denoising decoder parameterized by $\theta$. The noise prediction $\epsilon_{\theta}$ is conditioned on the noisy state $x_{t}$, the timestep $t$, and the feature group featfeat. The parameters $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\sigma_{t}^{2}=\beta_{t}$ are derived from a predefined noise scheduler $\beta_{t}$. Eq. 2 describes how to estimate a slightly less noisy state ($x_{t-1}$) from the current noisy state ($x_{t}$) using the model’s noise prediction ($\epsilon_{\theta}$). The decoder is trained to predict the noise added during the forward process. The optimization objective involves minimizing the difference between the actual sampled noise $\epsilon$ and predicted noise $\epsilon_{\theta}$, formulated as a gradient descent step on:
[…]
where $\epsilon$ is randomly sampled from $\mathcal{N}(0,I)$. Detailed computation of trajectory denoising loss, $\mathcal{L}_{diff}$, using Mean Squared Error (MSE) is outlined in Algorithm 1. In detail, TransDiffuser denoises an 8-step action sequence in parallel during inference which is then cumulatively summed to recover the multiple key waypoints of the complete trajectory.
Input: Trajectory GT labellabel, Scene feature featfeat, Number of denoising timesteps $\mathbf{T}$
@@
1 // Generate random Gaussian noise ;
2 $noise\leftarrow\text{Gaussian}(label.shape)$ ;
3 // Pick a random time step ;
4 $step\leftarrow\text{Rand}(0,\mathbf{T})$ ;
5 // Apply noise in the forward process
6 $noisy\_target\leftarrow\text{AddNoise}(label,noise,step)$ ;
7 // Denoising process
8 $pred\leftarrow\text{Denoising\_Decoder}(noisy\_target,step,feat)$ ;